Widmer Company had gross wages of $240,000 during the week ended Jun 17. The amount of wages subject to social security tax was $240,000, while the amount of wages subject to federal and state unemployment taxes was $35,000. Tax rates are as follows: Social Security 6% Medicare 1.5% State unemployment 5.4% Federal unemployment 0.8% The total amount withheld from employee wages for federal taxes was $48,000. a. Journalize the entry to record the payroll employees earning report for the week of June 17.

b: Journalize the entry to record the payroll tax expense incurred for the week of June 17 on Dec 27.
Answer #1
a
Wages expense 240000
     Social Security Tax payable 14400 =240000*6%
     Medicare Tax payable 3600 =240000*1.5%
     Employees Federal Income Tax payable 48000
    Wages payable 174000
b
Payroll tax expense 20170
     Social Security Tax payable 14400
     Medicare Tax payable 3600
     State unemployment tax payable 1890 =35000*5.4%
     Federal unemployment tax payable 280 =35000*0.8%